OpenStudy (anonymous):

HELP PLEASE! Factor Each Expression 54. -2h^2 + 4h + 70 = 68. w^2 + 24w + 144 = 70. 25x^2 - 36 =

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

exercise #70 we have to apply this identity: \[{a^2} - {b^2} = \left( {a - b} \right)\left( {a + b} \right)\]

OpenStudy (michele_laino):

where \(a^2=25, b^2=36\)

OpenStudy (michele_laino):

oops...\(a^2=25x^2\)

OpenStudy (anonymous):

70. 25x^2 - 36 = 25x^2 - 36 right?

OpenStudy (michele_laino):

please, we have to express the left side as a product between two factors

OpenStudy (anonymous):

Can you walk me threw it please? I don't understand what to do. You said a^2 - b^2 = (a - b) (a + b) So what do I do?

OpenStudy (michele_laino):

we have: \[25{x^2} = {\left( {5x} \right)^2}\] so if \(a^2=25x^2\) then \(a=5x\)

OpenStudy (anonymous):

Okay, so 70. 25x^2 - 36 = 25x^2 = (5x)^2 Now what?

OpenStudy (michele_laino):

no, I'm sorry, we can write this: \(36=6^2\) so \(b^2=36\) then \(b=6\)

OpenStudy (michele_laino):

substituting into my formula above, i get: \[25{x^2} - 36 = \left( {5x - 6} \right)\left( {5x + 6} \right)\]

OpenStudy (anonymous):

70. 25x^2 - 36 = 25x^2 - 36 = (5x - 6) (5x + 6) Thats the answer?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Can you help me with the other two please

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

54. -2h^2 + 4h + 70 = 68. w^2 + 24w + 144 =

OpenStudy (michele_laino):

what is \((w+12)^2=...?\)

OpenStudy (anonymous):

12w^2 ?

OpenStudy (michele_laino):

no, I'm sorry, it is a square of binomial, here is the rule: the square of the first term--->\(w^2\) the square of the second term--->\(12^2=144\) the double product of the first term with the second term--->\(2\cdot w\cdot 12=24w\) please add this three quantities, what do you get?

OpenStudy (michele_laino):

these*

OpenStudy (anonymous):

68. w^2 + 24w + 144 = 24w, 144, w^2 ?

OpenStudy (michele_laino):

yes! you have to sum them each other, so you will get this: \[{w^2} + 24w + 144 = {\left( {w + 12} \right)^2}\]

OpenStudy (anonymous):

Thank you. So the answer is (w + 12)^2

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

exercise #54 we can rewrite your expression as below: \[ - 2{h^2} + 4h + 70 = - 2\left( {{h^2} - 2h - 35} \right)\]

OpenStudy (michele_laino):

am I right?

OpenStudy (anonymous):

How do you get -2h + 4h + 70 = -2 (h^2 - 2h - 35)

OpenStudy (michele_laino):

I divide each term by \(-2\)

OpenStudy (michele_laino):

divided*

OpenStudy (michele_laino):

and I factored out \(-2\)

OpenStudy (anonymous):

Ok, so after you factored out -2, now what?

OpenStudy (michele_laino):

I got this: \[ - 2{h^2} + 4h + 70 = - 2\left( {{h^2} - 2h - 35} \right)\]

OpenStudy (anonymous):

And -2 (h^2 – 2h – 35) is the answer?

OpenStudy (michele_laino):

please, wait. Now, you have to try to search for two numbers, such that their sum is \(2\), and their product is \(-35\)

OpenStudy (michele_laino):

for example: \(-7\) and \(5\) can be the correct numbers?

OpenStudy (michele_laino):

they are not the correct numbers, right?

OpenStudy (anonymous):

it should be 7 + -5

OpenStudy (michele_laino):

right! the correct numbers are \(7\) and \(-5\)

OpenStudy (michele_laino):

so, we can write this:

OpenStudy (anonymous):

Okay what?

OpenStudy (michele_laino):

\[\begin{gathered} - 2{h^2} + 4h + 70 = - 2\left( {{h^2} - 2h - 35} \right) = \hfill \\ \hfill \\ = - 2\left( {h - 7} \right)\left( {h - \left( { - 5} \right)} \right) = - 2\left( {h - 7} \right)\left( {h + 5} \right) \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

Okay, and -2 (h - 7) (h + 5) is the answer?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

Thanks! Can you help me with a few more please?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

please open a new question